# SoftVQ-VAE: Efficient 1-Dimensional Continuous Tokenizer

### v1

@@ Abstract
Efficient image tokenization with high compression ratios remains a critical challenge for training generative models. We present SoftVQ-VAE, a continuous image tokenizer that leverages soft categorical posteriors to aggregate multiple codewords into each latent token, substantially increasing the representation capacity of the latent space. When applied to Transformer-based architectures, our approach compresses 256$\times$256 and 512$\times$512 images using as few as 32 or 64 1-dimensional tokens. Not only does SoftVQ-VAE show consistent and high-quality reconstruction, more importantly, it also achieves state-of-the-art and significantly faster image generation results across different denoising-based generative models. Remarkably, SoftVQ-VAE improves inference throughput by up to 18x for generating 256$\times$256 images and 55x for 512$\times$512 images while achieving competitive FID scores of 1.78 and 2.21 for SiT-XL. It also improves the training efficiency of the generative models by reducing the number of training iterations by 2.3x while maintaining comparable performance. With its fully-differentiable design and semantic-rich latent space, our experiment demonstrates that SoftVQ-VQE achieves efficient tokenization without compromising generation quality, paving the way for more efficient generative models. Code and model are released  [^1] https://github.com/Hhhhhhao/continuous_tokenizer..
@@ 2.2 Denoising-based Generative Models
SiT [72] instead uses stochastic interpolants [1] and performs denoising via probability flow ordinary differential equation (PF ODE). The model learns the ODE’s velocity field [65] by minimizing mean-squared-error (MSE) with the interpolated noise. This velocity field defines a score function for sample generation, analogous to DiT’s reverse-time SDE.
@@ 4.1 Experiments Setup
Implementation Details of SoftVQ-VAE. We use the LlamaGen codebase [98] to build our SoftVQ-VAE. We instantiate 4 configurations: SoftVQ-S, SoftVQ-B, SoftVQ-BL, SoftVQ-L, with a total of 45M, 173M, 391M, and 608M parameters, respectively. Each configuration has variants of latent codes $L=64$ and $L=32$. For the representation alignment, we choose DINOv2 [79] for the pre-trained features, similarly as Yu et al. [116] and Li et al. [62]. To learn a better semantic latent space, we initialize our encoder from the pre-trained DINOv2 weights. Alignment with other pre-trained features is explored in Sec. 4.4. We train the tokenizers on ImageNet [15] of resolution 256$\times$256 and 512$\times$512 for 250K iterations. Similarly to Tian et al. [99] and Li et al. [62], we found that the discriminator is very important for training tokenizer, and we adopted the same frozen DINO-S [9, 79] discriminator with a similar architecture to StyleGAN [48, 49]. In addition, we use DiffAug [121], consistency regularization [119], and LeCAM regularization [101] for discriminator training as in [99]. For the training objective, we set $\lambda_{1}=1.0$, $\lambda_{1}=0.2$, $\lambda_{3}=0.1$, and $\lambda_{4}=0.01$, following previous common practice. More training details are shown in Sec. C.1.
@@ 4.2 Main Results
Efficiency. The system-level comparison reveals SoftVQ’s exceptional efficiency across multiple metrics. When integrated with MAR-H at 256$\times$256 resolution, SoftVQ reduces GFLOPs by  40% compared to standard KL tokenizer, i.e., 86.55 vs 145.08 GFLOPs with 64 tokens and further reducing to 67.93 GFLOPs with 32 tokens. The efficiency gains are even more pronounced in diffusion-based models, especially with higher resolutions. SoftVQ enables DiT-XL and SiT-XL to achieve significantly lower computational costs i.e., 28.81 and 29.23 GFLOPs, respectively with 64 tokens, compared to their KL counterparts (80.73 GFLOPs with 1024 tokens). Importantly, these efficiency improvements come without compromising generation quality. SiT-XL trained on SoftVQ-BL with 64 tokens presents  55x throughput than the KL counterpart, while showing competitive IS and FID scores in all configurations. The improved GFLOPs also lead to a more efficient training. Not only do we require fewer training iterations to achieve a comparable performance to leading systems, i.e., 2.5M vs 7M steps of SiT-XL and DiT-XL, the high compression ratio of SoftVQ also improves the training throughput, reducing the time to train SiT-XL for 400K steps from 72 to 20 hours on 8 MI250.
@@ 4.3 Comparison of Tokenizers
Tokenizer # Params # Tokens Token Dim. rFID $\downarrow$ SiT-L
@@
Less Lossy Property. SoftVQ-S exhibits remarkably consistent performance across different compression ratios. When reducing tokens from 256 to 32, SoftVQ-S maintains a minimal degradation in rFID from 0.80 to 1.24, significantly outperforming both VQ-S, from 1.45 to 10.97, and AE-S, from 1.15 to 2.01. The benefits of the fully-differentiable property of SoftVQ along with representation alignment are particularly evident in generation quality, where SoftVQ-S achieves substantially better gFID of 9.21-12.89, and IS scores of 79.5-93.6, compared to baselines, showing its robustness in preserving information at high compression ratios and a better latent space for generative models to be trained on.
@@ D.2 Main Results
D.3 Reconstruction Results

### v3

@@ Abstract
Efficient image tokenization with high compression ratios remains a critical challenge for training generative models. We present SoftVQ-VAE, a continuous image tokenizer that leverages soft categorical posteriors to aggregate multiple codewords into each latent token, substantially increasing the representation capacity of the latent space. When applied to Transformer-based architectures, our approach compresses 256$\times$256 and 512$\times$512 images using as few as 32 or 64 1-dimensional tokens. Not only does SoftVQ-VAE show consistent and high-quality reconstruction, more importantly, it also achieves state-of-the-art and significantly faster image generation results across different denoising-based generative models. Remarkably, SoftVQ-VAE improves inference throughput by up to 18x for generating 256$\times$256 images and 55x for 512$\times$512 images while achieving competitive FID scores of 1.78 and 2.21 for SiT-XL. It also improves the training efficiency of the generative models by reducing the number of training iterations by 2.3x while maintaining comparable performance. With its fully-differentiable design and semantic-rich latent space, our experiment demonstrates that SoftVQ-VAE achieves efficient tokenization without compromising generation quality, paving the way for more efficient generative models. Code and model are released  [^1] https://github.com/Hhhhhhao/continuous_tokenizer..
@@ 2.2 Denoising-based Generative Models
SiT [72] instead uses stochastic interpolants [1] and performs denoising via probability flow ordinary differential equation (PF ODE). The model learns the ODE’s velocity field [65] by minimizing mean-squared-error (MSE) loss. This velocity field defines a score function for generation, analogous to DiT’s reverse-time SDE.
@@ 4.1 Experiments Setup
Implementation Details of SoftVQ-VAE. We use the LlamaGen codebase [98] to build our SoftVQ-VAE. We instantiate 4 configurations: SoftVQ-S, SoftVQ-B, SoftVQ-BL, SoftVQ-L, with a total of 45M, 173M, 391M, and 608M parameters, respectively. Each configuration has variants of latent codes $L=64$ and $L=32$. For the representation alignment, we choose DINOv2 [79] for the pre-trained features, similarly as Yu et al. [116] and Li et al. [62]. To learn a better semantic latent space, we initialize our encoder from the pre-trained DINOv2 weights. Alignment with other pre-trained features is explored in Sec. 4.4. We train the tokenizers on ImageNet [15] of resolution 256$\times$256 and 512$\times$512 for 250K iterations. Similarly to Tian et al. [99] and Li et al. [62], we adopted the same frozen DINO-S [9, 79] discriminator with a similar architecture to StyleGAN [48, 49]. We use DiffAug [121], consistency regularization [119], and LeCAM regularization [101] for discriminator training as in [99]. For the training objective, we set $\lambda_{1}=1.0$, $\lambda_{1}=0.2$, $\lambda_{3}=0.1$, and $\lambda_{4}=0.01$, following previous common practice. More training details are shown in Sec. C.1.
@@ 4.2 Main Results
Efficiency. The system-level comparison reveals SoftVQ’s exceptional efficiency across multiple metrics. When integrated with MAR-H at 256$\times$256 resolution, SoftVQ reduces GFLOPs by  40% compared to standard KL tokenizer, i.e., 86.55 vs 145.08 GFLOPs with 64 tokens and further reducing to 67.93 GFLOPs with 32 tokens. SoftVQ enables DiT-XL and SiT-XL to achieve significantly lower computational costs i.e., 28.81 and 29.23 GFLOPs, respectively with 64 tokens, compared to their KL counterparts (80.73 GFLOPs with 1024 tokens). Importantly, these efficiency improvements come without compromising generation quality. SiT-XL trained on SoftVQ-BL with 64 tokens presents  55x throughput than the KL counterpart, while showing competitive IS and FID scores in all configurations. Not only do we require fewer training iterations to achieve a comparable performance to leading systems, i.e., 2.5M vs 7M steps of SiT-XL and DiT-XL, the high compression ratio of SoftVQ also improves the training throughput, reducing the time to train SiT-XL for 400K steps from 72 to 20 hours on 8 MI250.
@@ 4.3 Comparison of Tokenizers
Tokenizer # Params # Tokens Dim. rFID $\downarrow$ SiT-L
@@
Less Lossy Property. SoftVQ-S exhibits remarkably consistent performance across different compression ratios. When reducing tokens from 256 to 32, SoftVQ-S maintains a minimal degradation in rFID from 0.80 to 1.24, significantly outperforming both VQ-S, from 1.45 to 10.97, and AE-S, from 1.15 to 2.01. The benefits of the fully-differentiable property of SoftVQ along with representation alignment are particularly evident in generation quality, where SoftVQ-S achieves substantially better gFID of 9.21-12.89, and IS scores of 79.5-93.6, compared to baselines, showing its robustness in preserving information at high compression ratios and a better latent space for generative models.
@@ Appendix A Posterior of KL-VAE and VQ-VAE
A.1 Comparison to VQ
We list the advantages of SoftVQ in Tab. 5 with empirical support, and will add in the revision.
Table 5: Advantages of SoftVQ over VQ.
Advantages VQ SoftVQ Empirical Support
Grad. Broken yes no
No straight-through trick. SoftVQ-S-64 has a 1.41 rFID and 13.35 gFID.
Codebook Loss yes no
Not necessary. Adding codebook loss in SoftVQ-S-64
results in an 1.34 rFID of of but a worse gFID of 15.15.
Commit Loss yes no
Not necessary. Adding commit loss in SoftVQ-S-64
results in an rFID of 1.33 of but a worse gFID of 14.35.
Comp. Ratio low high
SoftVQ has much lower rFID$<$1.0 comapred to TiTok in Tab. 1.
Rep. Align. difficult easy
More discriminative features in main Fig. 4
While representation alignment can also applied to VQ, it fails to learn discriminative features with high compression ratio, mainly due to its broken gradient. Compared to SoftVQ, whose encoder and codebook’s parameters are directly learned by the alignment loss, the gradient of alignment loss on VQ will first be straight-through from decoder input to encoder output, and then the codebook is updated according to the codebook loss, resulting in indirect learning of discriminative features.
We provide an additional codewords visualization in Fig. 5, where SoftVQ learns code embeddings uniformly across the entire distribution.
Figure 5: Codewords visualization
@@ D.2 Main Results
D.3 More Comparison to TiTok
We provide a comparison between TiTok and SoftVQ in Tab. 10, including token numbers, model parameters, rFID of IN-1K and MSCOCO, linear probing (LP) accuracy, gFID and IS of SiT-L. We show that SoftVQ significantly outperforms TiTok with VQ and KL. Also, we need to note that while the model architectures are similar, training is different. TiTok is trained using 2-stage decoders with a pre-trained tokenizer, whereas our method is trained end-to-end with a single decoder. Also note that, although soft assignment of codewords may not be a new idea, our work is indeed the first to explore it with continuous tokenizer that achieves high compression ratio.
Table 10: Comparison between TiTok and SoftVQ.
Tokenizer # Param. rFID (IN-1K) rFID (COCO) LP (IN-1K) gFID IS
TiTok-B-64 176 1.70 11.20 39.2 25.12 53.1
TiTok-BL-KL-64 389 1.25 8.94 11.4 23.35 52.7
TiTok-BL-VQ-64 390 2.60 11.24 9.3 19.23 61.8
SoftVQ-B-64 173 0.89 5.16 42.3 10.13 103.4
TiTok-L-32 614 2.21 14.41 48.9 26.84 49.5
TiTok-LL-KL-32 614 1.61 10.45 12.3 24.65 51.97
SoftVQ-L-32 608 0.61 5.51 59.4 9.47 107.2
D.4 Reconstruction Visualization